40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Feb 10 18:00:13 2013, yuta, Summary, LSC, PR2-flipped half-PRC spectra/OLTF POPDCLSCPRM.pngLSCPRCLOLTF.png
    Reply  Mon Feb 11 12:49:54 2013, Jenne, Summary, LSC, Resonant freq change - why? (and passive TT mode freqs) 
Message ID: 8054     Entry time: Mon Feb 11 12:49:54 2013     In reply to: 8053
Author: Jenne 
Type: Summary 
Category: LSC 
Subject: Resonant freq change - why? (and passive TT mode freqs) 

Quote:

  Is it because of the change in the resonant frequency of the BS-PRM stack? How much the load on BS-PRM changed?
  Or is it because of the change in the resonant frequency of PR2/PR3

I claim that neither of those things is plausible.  We took out 1 PZT, and put in 1 active TT onto the BS table.  There is no way the resonant frequency changed by an appreciable amount due to that switch.

I don't think that it is the resonant frequency of the TTs either.  Here, I collate the data that we have on the resonant frequencies of our tip tilts.  It appears that in elog 3425 I recorded results for TTs 2 and 3, but in elog 3447 I just noted that the measurements had been done, and never put them into the elog.  Ooops.

Resonant frequency and Q of modes of passive tip tilts. 

  Vertical Yaw Pos Side
TT1 f0=20, Q=18 f0=1.89, Q=3.8 f0=1.85, Q=2 f0=1.75, Q=3.2
TT2 f0=24, Q=7.8 f0=1.89, Q=2.2 f0=1.75, no Q meas f0=1.8, Q=4.5
TT3 f0=20, Q=34 f0=1.96, Q="low" f0=1.72, Q=3.3 f0=1.85, Q=6
TT4 f0=21, Q=14 f0=1.88, Q=2.3 f0=1.72, Q=1.4 f0=1.85, Q=1.9
TT5 f0=20, Q=22.7 no measurement f0=1.79, Q=1.8 f0=1.78, Q=3.5

Notes:  "Serial Number" of TTs here is based on the SN of the top suspension point block.  This does not give info about which TT is where.  Pitch modes were all too low of Q to be measured, although we tried.

Tip tilt mode measurements were taken with a HeNe and PD shadow sensor setup - the TT's optic holder ring was partially obscuring the beam.

ELOG V3.1.3-